### v1

# Is Repeated Bayesian Interim Analysis Consequence-Free?

Abstract
Interim analyses are commonly used in clinical trials to enable early stopping for efficacy, futility, or safety. While their impact on frequentist operating characteristics is well studied and broadly understood, the effect of repeated Bayesian interim analyses—when conducted without appropriate multiplicity adjustment— remains an area of active debate. In this article, we provide both theoretical justification and numerical evidence illustrating how such analyses affect key inferential properties, including bias, mean squared error, the coverage probability of posterior credible intervals, false discovery rate, familywise error rate, and power. Our findings demonstrate that Bayesian interim analyses can significantly alter a trial’s operating characteristics, even when the prior used for Bayesian inference is correctly specified and aligned with the data-generating process. Extensive simulation studies, covering a variety of endpoints, trial designs (single-arm and two-arm randomized controlled trials), and scenarios with both correctly specified and misspecified priors, support theoretical insights. Collectively, these results underscore the necessity of appropriate adjustment, thoughtful prior specification, and comprehensive evaluation to ensure valid and reliable inference in Bayesian adaptive trial designs.
KEY WORDS: Bayesian adaptive design; interim analysis; interim stopping; bias; coverage probability; false discovery rate.
1 Introduction
In clinical trials, interim analyses are often conducted to assess trial progress and inform decisions about whether to continue or stop the trial early. A trial may be stopped early if interim results cross predefined thresholds, such as showing compelling efficacy, indicating futility due to lack of benefit, or revealing unacceptable toxicity that compromises participant safety. Two primary types of decision rules are used in interim analyses and stopping: frequentist and Bayesian.
In the frequentist approach, the impact of interim analyses on the operating characteristics of clinical trials—particularly the familywise error rate (FWER)—has been extensively studied and is generally understood. Within this framework, parameters are treated as fixed but unknown constants. Repeated interim looks introduce a multiplicity problem, leading to inflation of the FWER. To address this, various methods have been developed to adjust significance levels or decision rules at each interim analysis, thereby ensuring that the overall FWER remains controlled at the desired level. The Pocock method (Pocock, 1977) assigns equally stringent significance levels to all interim analyses, making early stopping easier but potentially sacrificing efficiency and power. In contrast, the O’Brien-Fleming method (1979) applies more stringent stopping criteria early in the trial and gradually relaxes them as the trial progresses, offering a more conservative approach. Alpha spending functions (DeMets and Lan, 1983) provide additional flexibility by allocating the total alpha level across multiple analyses and dynamically adjusting critical values based on how much alpha has already been ‘spent’.
In contrast, the Bayesian framework—which treats parameters as random variables with associated probability distributions—has been less extensively studied in the context of interim analyses, and there is no clear consensus on their impact on trial operating characteristics. Berry (1987) argued that Bayesian interim analyses are not subject to multiplicity issues, as they adhere to the likelihood principle, allowing analyses to be conducted at any point during the trial without requiring adjustments. Harrell (2017) echoed this perspective, advocating the use of Bayesian posterior probabilities to make go/no-go decisions as frequently as needed. However, this view has not gone unchallenged. Jennison and Turnbull (1990) cautioned that Bayesian methods may exhibit poor frequentist properties, including an inflated FWER when multiple interim looks are performed.
Beyond FWER control, less attention has been given to other statistical properties, such as bias, mean square error (MSE), and interval coverage in treatment effect estimation. In the frequentist framework, Armitage (1958) pointed out that multiple interim looks can impact the maximum likelihood estimator and the coverage probability of the confidence interval. Tsiatis, Rosner, and Mehta (1984) described a method of constructing confidence intervals with exact coverage probabilities for the mean of a normal distribution. In the Bayesian framework, Rosenbaum and Rubin (1984) conducted simulation studies to examine the coverage probability of the posterior credible interval under a flat user-specified prior, when the data-generating prior was a normal distribution with varying precision. They found that coverage probabilities deviated from nominal value when repeated Bayesian interim analyses were applied, particularly when the precision of the data-generating prior increased. More recently, Bunouf (2023) proposed the Beta-J prior for single-arm binomial trials, demonstrating improved frequentist properties, such as reduced bias and enhanced interval coverage, compared to standard conjugate Beta priors.
Despite these findings, most research has prioritized FWER control over evaluating other critical inferential quantities, such as bias, MSE, and interval coverage. Recent work by Robertson et al. (2023) and Bunouf (2023) has highlighted this gap, calling for more comprehensive assessments of how early trial termination affects treatment effect estimates. The U.S. Food and Drug Administration has also acknowledged this concern, noting in recent guidance (FDA, 2019) that, compared to the FWER inflation, “biased estimation in adaptive design is currently a less well-studied phenomenon.”
In this article, we aim to address this critical knowledge gap by providing both theoretical and numerical examinations of how repeated Bayesian interim analyses affect key statistical properties of clinical trial designs. Specifically, we assess their impact on bias, MSE, coverage probability of the posterior credible intervals, false discovery rate (FDR), FWER, and power. Our analysis explicitly distinguishes between the data-generating prior and the user-specified prior. We show that even when the two priors are identical, repeated Bayesian interim analyses can systematically distort some key inferential quantities. These distortions become even more pronounced when the user-specified prior is misspecified. To support our theoretical findings, we conduct extensive simulation studies across a wide range of settings, including different types of endpoints (binary, continuous, and time-to-event) and both single-arm and two-arm randomized controlled trials (RCTs), highlighting the generality and practical significance of our results.
The remainder of this article is organized as follows. In Section 2, we present Bayesian probability models for binary, normal, and time-to-event outcomes, along with a theoretical assessment of the impact of repeated Bayesian interim analyses on the operating characteristics of clinical trials. Section 3 evaluates these effects through extensive simulation studies. Finally, we offer concluding remarks in Section 4.
2 Methods
2.1 Probability Models
Consider a superiority trial that aims to demonstrate the superiority of an experimental treatment over a control, with interim analyses that may lead to early termination if there is overwhelming evidence of superiority. Let $X_{i}$ denote the efficacy endpoint for the $i$th patient receiving the experimental treatment, where $X_{i}$ follows a probability distribution indexed by the parameter $\theta$, representing the treatment effect of interest, possibly along with other nuisance parameters. Without loss of generality, we assume that larger values of $\theta$ indicate higher efficacy. The hypotheses of interest are formulated as:
$H_{0}:\theta-\theta_{0}\leq\delta\quad\mathrm{vs.}\quad H_{a}:\theta-\theta_{0}>\delta,$
where $\theta_{0}$ denotes the treatment effect for the control, and $\delta$ is an equivalence margin, which is often set at 0 in superiority trials, and a negative value for non-inferiority trials. In single-arm trials, $\theta_{0}$ is assumed to be a known constant, while in two-arm RCTs, $\theta_{0}$ is treated as an unknown parameter. While we adopt this formulation for clarity of exposition, alternative hypothesis structures, such as those based on hazard ratios for time-to-event endpoints in RCTs, are also possible and are discussed later.
In the Bayesian framework, the parameter $\theta$ is treated as a random variable with an associated probability distribution. For our analysis, it is essential to distinguish between two types of priors for $\theta$: the data-generating prior and the user-specified prior. The data-generating prior distribution of $\theta$ represents the true underlying distribution that generates $\theta$, which in turn generates the observed data $X_{i}$. In practice, the data-generating prior is typically unknown. The user-specified prior, on the other hand, refers to the prior assumed by users for Bayesian inference. It can be constructed based on prior knowledge, historical data, or chosen to be noninformative or vague to limit its impact on the posterior. In nearly all applications, the user-specified prior is misspecified in that it differs from the true data-generating prior.
Next, we present commonly used Bayesian models for binary, normal, and time-to-event endpoints, along with their associated prior and posterior distributions. For binary and normal outcomes in RCTs, patients in the treatment and control arms are independent, allowing the same model to be applied separately to each arm. As a result, we do not describe separate models for single-arm trials and RCTs in these cases. For time-to-event endpoints, considering the popularity of using the Cox proportional hazards model (Cox, 1972) in RCTs, we present separate modeling approaches for single-arm trials and RCTs. These models are not required for the theoretical development presented in the next section, but they facilitate the understanding of theoretical results and serve as the foundation for the simulation study.
Binary endpoint: For a binary endpoint $X_{i}$, such as tumor response, we assume
$X_{i}\sim\mbox{Bernoulli}(\theta),$ (1)
where $\theta$ denotes the response rate. Suppose the user-specified prior for $\theta$ is $\pi(\theta)=\mbox{Beta}(\alpha,\beta)$, where $\alpha$ and $\beta$ are hyperparameters. Given interim data ${\cal D}_{n}=(X_{1},\cdots,X_{n})$ from the first $n$ patients in the trial, the posterior distribution for $\theta$ under the user-specified prior is
$\pi(\theta|{\cal D}_{n})=\mbox{Beta}(\alpha+\sum_{i=1}^{n}X_{i},\beta+n-\sum_{i=1}^{n}X_{i})$
with posterior mean $\hat{\theta}=\frac{\alpha+\sum_{i=1}^{n}X_{i}}{\alpha+\beta+n}$.
Normal endpoint: For a continuous outcome, we assume
$X_{i}\sim N(\theta,\sigma^{2}),$ (2)
where $\theta$ and $\sigma^{2}$ represent the mean and variance, respectively. Suppose the user-specified joint prior distribution of $\theta$ and $\sigma^{2}$ is a conjugate prior $\pi(\theta,\sigma^{2})=\pi(\theta|\sigma^{2})\pi(\sigma^{2})$, with $\pi(\theta|\sigma^{2})=N(\mu,\sigma^{2}/\kappa)$ and $\pi(\sigma^{2})=\mbox{Inv-}\chi^{2}(\nu,\sigma_{0}^{2})$, where $\mu$, $\kappa$, $\nu$, and $\sigma_{0}^{2}$ are hyperparameters. Here, $\nu$ and $\sigma_{0}^{2}$ denote the degrees of freedom and scale parameter for the inverse chi-squared distribution. Given interim data ${\cal D}_{n}=(X_{1},\cdots,X_{n})$ from the first $n$ patients with sample mean $\bar{X}_{n}=(X_{1}+\cdots+X_{n})/{n}$, the marginal posterior distribution of $\theta$ under the user-specified prior is
$\pi(\theta|{\cal D}_{n})=t_{\nu+n}(\nu_{n},\tau_{n}),$
where the location (also the posterior mean, if $\nu+n>1$) and scale parameters are given by
$\nu_{n}=\frac{n\bar{X}_{n}+\kappa\mu}{\kappa+n}$
$\tau_{n}=\big{\{}\nu\sigma_{0}^{2}+\sum(X_{i}-\bar{X}_{n})^{2}+\frac{\kappa n}{\kappa+n}(\bar{X}_{n}-\mu)^{2}\big{\}}/(\nu+n)(\kappa+n).$
In order to focus on the effect of the prior distribution of $\theta$ on interim analyses, we also consider the case where the variance $\sigma^{2}$ is known. Specifically, we assume the data $X_{i}$ is standardized and set $\sigma^{2}=1$. In this setting, the user-specified prior for $\theta$ is the conjugate prior $\theta\sim N(\mu,\sigma_{0}^{2})$, with $\mu$ and $\sigma_{0}^{2}$ as hyperparameters. Given interim data ${\cal D}_{n}$ with sample mean $\bar{X}_{n}$, the posterior distribution for $\theta$ under the user-specified prior is $\pi(\theta|{\cal D}_{n})=N(\mu_{n},\sigma_{n}^{2})$, where the posterior mean $\mu_{n}$ and variance $\sigma_{n}^{2}$ are given by
$\mu_{n}=\frac{\frac{1}{\sigma_{0}^{2}}\mu+\frac{n}{\sigma^{2}}\bar{X}_{n}}{\frac{1}{\sigma_{0}^{2}}+\frac{n}{\sigma^{2}}}\ \ \ \ \ \ \mbox{and}\ \ \ \ \ \ \frac{1}{\sigma_{n}^{2}}=\frac{1}{\sigma_{0}^{2}}+\frac{n}{\sigma^{2}}.$ (3)
Single-arm trials with time-to-event endpoint: For a time-to-event endpoint, such as overall survival or progression-free survival, we assume that $X_{i}$ follows a Weibull distribution parameterized in terms of the median $\theta$ and the shape parameter $\kappa$:
$X_{i}\sim\mbox{Weibull}(\theta,\kappa),$ (4)
with probability density function
$f(X_{i}|\theta,\kappa)=(ln2)\kappa\theta^{-\kappa}{X_{i}}^{\kappa-1}\mbox{exp}\big{\{}-(ln2)\theta^{-\kappa}{X_{i}}^{\kappa}\big{\}},$ (5)
and survival function
$S(X_{i}|\theta,\kappa)=\mbox{exp}\big{\{}-ln(2)\theta^{-\kappa}X_{i}^{\kappa}\big{\}}.$ (6)
This parameterization expresses the Weibull distribution in terms of the median $\theta$, which often has a more interpretable clinical interpretation, and the shape parameter $\kappa$. At any interim decision time, let $t_{i}$ denote the observed time to event or censoring time for the $i$th patient, and let $\zeta_{i}=I(t_{i}=X_{i})$ denote the censoring indicator. The likelihood for patient $i$, based on data $(t_{i},\zeta_{i})$, is given by $L\big{(}(t_{i},\zeta_{i})|\theta,\kappa\big{)}=f(t_{i}|\theta,\kappa)^{\zeta_{i}}S(t_{i}|\theta,\kappa)^{1-\zeta_{i}}$. Given interim data ${\cal D}_{n}=\big{(}(t_{1},\zeta_{1}),\cdots,(t_{n},\zeta_{n})\big{)}$ from the first $n$ patients, the joint likelihood function is $L({\cal D}_{n}|\theta,\kappa)=\prod_{i=1}^{n}L\big{(}(t_{i},\zeta_{i})|\theta,\kappa\big{)}$.
We assume user-specified priors $\pi(\theta)=\mbox{Gamma}(\phi,\eta)$ and $\pi(\kappa)=\mbox{Gamma}(\lambda,\gamma)$ for the median $\theta$ and shape parameter $\kappa$, respectively. Given interim data ${\cal D}_{n}$, the joint posterior distribution of $(\theta,\kappa)$ under the user-specified prior is $\pi(\theta,\kappa|{\cal D}_{n})\propto\pi(\theta)\pi(\kappa)L({\cal D}_{n}|\theta,\kappa)$, which can be sampled using the Gibbs sampler (Geman and Geman, 1984; Gelfand and Smith, 1990).
RCT with time-to-event outcomes: For an RCT with time-to-event outcomes, we assume a Cox proportional hazards model of the form
$h(t|\theta,\kappa,\beta)=h_{0}(t|\theta,\kappa)\mbox{exp}(\beta Z),$ (7)
where $Z$ is the treatment indicator ($Z=1$ for the experimental treatment and $Z=0$ for the control), and $\mbox{exp}(\beta)$ represents the hazard ratio (HR) comparing the experimental treatment to the control. We consider the following hypotheses test: $H_{0}:\mbox{exp}(\beta)>\rho$ versus $H_{a}:\mbox{exp}(\beta)<\rho$, where $\rho$ is a predefined upper bound for HR used to define superiority.
We specify the baseline hazard function $h_{0}(t|\theta,\kappa)=(ln2))\theta^{-\kappa}\kappa t^{\kappa-1}$, corresponding to a Weibull distribution parameterized by the median $\theta$ and shape parameter $\kappa$. The resulting probability density function and survival function are
$f(t|\theta,\kappa,\beta)=(ln2)\kappa\theta^{-\kappa}{t}^{\kappa-1}\mbox{exp}(\beta Z)\mbox{exp}(-(ln2)\theta^{-\kappa}\mbox{exp}(\beta Z){t}^{\kappa}),$ (8)
and
$S(t|\theta,\kappa,\beta)=\mbox{exp}\big{(}-ln(2)\theta^{-\kappa}\mbox{exp}(\beta Z)t^{\kappa}\big{)}.$ (9)
The likelihood for patient $i$ with data $(t_{i},\zeta_{i})$ is $L\big{(}(t_{i},\zeta_{i})|\theta,\kappa,\beta\big{)}=f(t_{i}|\theta,\kappa,\beta)^{\zeta_{i}}S(t_{i}|\theta,\kappa,\beta)^{1-\zeta_{i}}$. Given interim data ${\cal D}_{n}=\big{(}(t_{1},\zeta_{1}),\cdots,(t_{n},\zeta_{n})\big{)}$ from the first $n$ patients, the joint likelihood function is $L({\cal D}_{n}|\theta,\kappa,\beta)=\prod_{i=1}^{n}L\big{(}(t_{i},\zeta_{i})|\theta,\kappa,\beta\big{)}$. We assume user-specified priors $\pi(\theta)=\mbox{Gamma}(\phi,\eta)$, $\pi(\kappa)=\mbox{Gamma}(\lambda,\gamma)$, and $\pi(\beta)=\mbox{Normal}(\xi,\tau^{2})$ for the parameters $\theta$, $\kappa$, and $\beta$, respectively. The joint posterior distribution of $(\theta,\kappa,\beta)$ under the user-specified prior is $\pi(\theta,\kappa,\beta|{\cal D}_{n})\propto\pi(\theta)\pi(\kappa)\pi(\beta)L({\cal D}_{n}|\theta,\kappa,\beta)$, which can be sampled using the Gibbs sampler.
2.2 Decision rule and inferential metrics
Let $N$ denote the maximum sample size of the trial and $n$ an interim sample size. We consider the following Bayesian superiority stopping rule:
Stop the trial and reject $H_{0}$ if $\mbox{Pr}(H_{a}|{\cal D}_{n})>C$, where $C\in(0,1)$ is a probability cutoff.
This decision rule is evaluated at each interim analysis, as well as at the end of the trial. Of note, throughout this article, we use $H_{a}$ as a plug-in term to literally represent $\theta-\theta_{0}>\delta$ (i.e., $\mbox{Pr}(H_{a}|{\cal D}_{n})\equiv\mbox{Pr}(\theta-\theta_{0}>\delta|{\cal D}_{n})$). For ease of exposition, we refer to the trial design with a fixed sample size $N$ and no interim analyses as the fixed design, and the design that incorporates interim analyses and allows early stopping as the adaptive design.
To isolate the effect of interim analyses, we assume that the cutoff $C$ used in the decision rule is pre-determined based on specific criteria, such as controlling FDR or type I error rate under the fixed design. This same cutoff is applied when evaluating the operating characteristics of both the fixed and adaptive designs, ensuring that any observed differences can be attributed solely to the impact of interim analyses.
To evaluate the impact of Bayesian interim analysis, we adhere to the fundamental assumption of the Bayesian framework that $\theta$ is random. This avoids the philosophical argument that evaluating Bayesian designs using the frequentist criteria (i.e., assuming $\theta$ is fixed, such as evaluating type I error at a fixed null value) is neither meaningful nor interpretable. This argument has been a main barrier to reaching a consensus on whether Bayesian methods can accommodate multiple interim looks without any consequences.
Treating $\theta$ as a random variable necessitates distinguishing between the data-generating prior $\pi_{g}(\theta)$ and the user-specified prior $\pi(\theta)$ used for Bayesian inference. We assume that the data are generated according to the Bayesian process: sampling $\theta$ from $\pi_{g}(\theta)$, and generating data ${\cal D}_{n}$ based on $f(X_{i}|\theta)$. $\pi_{g}(\theta)$ can be interpreted as the distribution of the treatment effect $\theta$ across the population of all trials within a disease area, including positive and negative trials, which is generally unknown. In practice, for a specific trial, users have to specify a prior to make posterior inferences and interim decisions. They assume a user-specified prior $\pi(\theta)$ (e.g., a noninformative prior) and, conditional on the observed data ${\cal D}_{n}$, calculate $\mbox{Pr}(\theta|{\cal D}_{n})$ to make go/no-go decisions.
We explore scenarios in which $\pi(\theta)$ and $\pi_{g}(\theta)$ are either identical or different; in this paper, we refer to these cases as $\pi(\theta)$ being correctly specified or misspecified, respectively. We assume that the data model $f(X_{i}|\theta)$ is always correctly specified to focus on the impact of interim analyses. To evaluate the impact of interim analyses on trial operating characteristics, we consider the following inferential quantities:
  • Bias of posterior mean: The expected difference between $\hat{\theta}=\mathbb{E}(\theta|{\cal D})$ and the true value of $\theta$, i.e., $\mathbb{E}_{\theta\sim\pi_{g},\mathcal{D}\sim f(\cdot\mid\theta)}[\hat{\theta}-\theta]$.
  • MSE of posterior mean: Defined as $\mathbb{E}_{\theta\sim\pi_{g},\mathcal{D}\sim f(\cdot\mid\theta)}[(\hat{\theta}-\theta)^{2}]$.
  • Coverage of 95% posterior credible intervals: The probability that a 95% posterior credible interval of $\theta$ contains the true value of $\theta$, i.e., $\Pr_{\theta\sim\pi_{g},\mathcal{D}\sim f(\cdot\mid\theta)}(\theta\in CI)$, where $CI$ represents a 95% credible interval for $\theta$ derived from the posterior distribution $\pi(\theta\mid\mathcal{D})$ under $\pi(\theta)$. We consider both one-sided and symmetric credible intervals, defined as $(\theta_{0.05},\infty)$ and $(\theta_{0.025},\theta_{0.975})$, respectively, where $\theta_{\alpha}$ represents the $\alpha$-th percentile of the posterior distribution $\pi(\theta\mid\mathcal{D})$. The one-sided posterior credible interval is considered here to match the hypotheses of interest, $H_{0}:\theta-\theta_{0}<\delta$ vs $H_{a}:\theta-\theta_{0}>\delta$.
[…]
  • Type I error rate A: The probability of rejecting $H_{0}$ when $H_{0}$ is true, where rejection of $H_{0}$ occurs if $\Pr(H_{a}|\mathcal{D})>C$.
  • Type I error rate B: The probability of rejecting $H_{0}$ when $H_{0}$ is true, where rejection of $H_{0}$ occurs if the one-sided posterior credible interval for $\theta$ excludes the null region (i.e., $(-\infty,\theta_{0}+\delta]$). That is, $H_{0}$ is rejected when there is no overlap between the credible interval and the null region defined by $H_{0}$.
  • Power: The probability of rejecting $H_{0}$ when $H_{0}$ is false, i.e., $\Pr(\text{Reject }H_{0}\mid\theta-\theta_{0}>\delta)$.
  • Mean sample size of the trial: The average number of patients enrolled until stopping (early or at maximum sample size).
A few remarks are warranted. First, in the definition and calculation of the above inferential quantities, $\theta$ is random. For example, when calculating Type I error rate A, $\theta$ is a random variable under $H_{0}$ (i.e., following $\pi_{g}(\theta)$ constrained by $\theta-\theta_{0}<\delta$), rather than being fixed at a point null as in the classical frequentist definition of Type I errors. This setup aims to preempt the argument that the frequentist Type I error is not meaningful within the Bayesian framework. Second, rather than relying solely on the Type I error rate, we also consider the pFDR and FDR, as one criticism of the Type I error rate is that it is conditional on the truth, which is unknown. What matters more in practice is, in positive decisions (e.g., approving a drug), how many decisions are truly positive (e.g., the approved drug is actually effective). From this perspective, the pFDR can be interpreted as the risk of regulatory regret. For this reason, in our simulation study, we choose to calibrate the cutoff $C$ using the pFDR instead of the Type I error rate. However, the specific criterion used to calibrate $C$ is not critical for our purposes, as the focus here is on evaluating the impact of incorporating interim stopping, relative to the fixed design without interim analyses. Lastly, while we use the generic notation $\theta$, in the context of RCTs with time-to-event outcomes, the bias, MSE, and coverage metrics are computed with respect to the HR, $\exp({\beta})$.
2.3 Impact of interim stopping
In this section, we examine the impact of incorporating interim analyses and early stopping on various inferential quantities, using the corresponding fixed design (with no interims) as the reference. Bayesian inference for the fixed design is well established, and using it as the reference avoids unnecessary debate over the frequentist versus Bayesian framework. This allows us to focus on whether incorporating interim analyses and stopping rules alters the operating characteristics of Bayesian inference—analogous to how interim analyses introduce multiplicity issues in the frequentist framework. Table 1 summarizes our findings, which are general and not specific to any particular model or the Bayesian decision rule discussed earlier. For example, the Bayesian decision rule previously discussed uses the same cutoff $C$ at both interim and final analyses. However, the results presented in this section also apply to scenarios where different cutoffs are used at the interim and final analyses, provided that the same final analysis cutoff is used for both the fixed and adaptive designs. In this way, the only distinction between the two designs lies in the inclusion of interim analyses.
Table 1: Impact of interim superiority stopping on inferential quantities, with respect to the fixed design without interims, assuming the same decision rule for rejecting $H_{0}$.
Quantity Interim Stopping ($\pi=\pi_{g}$) Justification Interim Stopping ($\pi\neq\pi_{g}$) Justification
Posterior mean bias Unbiased Follows from the law of total expectation under the correct prior. Biased Prior misspecification combined with selective stopping distorts the posterior mean.
Credible interval coverage Maintains nominal level Credible intervals are calibrated under the prior predictive distribution. Degraded (over or undercoverage) Misspecified prior yields miscalibrated posteriors.
Posterior mean MSE Increased Early stopping raises variance due to smaller sample size. Often further increased Both bias (from prior) and variance (from stopping) contribute to MSE inflation.
FDR/pFDR Increased Multiple interim looks increase the chance of stopping on spurious signals under $H_{0}$. Often further increased Prior misspecification distorts posterior probabilities, often worsening FDR.
Type I error rate A Increased More looks at the data raise the probability of false positives. Often further increased Same mechanism, compounded by posterior miscalibration.
Type I error rate B Distorted (inflated or deflated) As its definition is conditional on $H_{0}$, $\pi_{g}$ is effectively restricted to the parameter space under $H_{0}$, leading to mismach with $\pi$. Distorted (inflated or deflated) Same mechanism, compounded by posterior miscalibration.
Not necessarily increased, as it depends on whether $\pi$ is more optimistic or pessimistic relative to $\pi_{g}$ and $H_{0}$.
Bias of the posterior mean
Let $\mathcal{D}_{n}$ be the observed data at an interim stopping time $n<N$, governed by a pre-specified stopping rule (e.g., stop when $\mbox{Pr}(H_{a}|{\cal D}_{n})>C$), and let $\hat{\theta}_{n}=\mathbb{E}_{\pi}[\theta\mid\mathcal{D}_{n}]$ denote the posterior mean estimates based on user-specified prior $\pi(\theta)$. Under the data-generating model $\pi_{g}(\theta)$ and $f(\cdot\mid\theta)$, the bias of $\hat{\theta}_{n}$ is given by
$\text{Bias}(\hat{\theta}_{n})$ $=$ $\mathbb{E}_{\theta\sim\pi_{g},\mathcal{D}_{n}\sim f(\cdot\mid\theta)}[\hat{\theta}_{n}]-\mathbb{E}_{\pi_{g}}[\theta]$
$=$ $\mathbb{E}_{\pi_{g}}\left[\mathbb{E}_{\mathcal{D}_{n}\mid\theta}\left[\mathbb{E}_{\pi}[\theta\mid\mathcal{D}_{n}]\right]\right]-\mathbb{E}_{\pi_{g}}[\theta].$
When $\pi=\pi_{g}$, by the law of total expectation (Casella and Berger, 2002), we have
$\mathbb{E}_{\pi_{g}}\left[\mathbb{E}_{\mathcal{D}_{n}\mid\theta}\left[\mathbb{E}_{\pi}[\theta\mid\mathcal{D}_{n}]\right]\right]=\mathbb{E}_{\pi_{g}}\left[\mathbb{E}_{\mathcal{D}_{n}\mid\theta}\left[\mathbb{E}_{\pi_{g}}[\theta\mid\mathcal{D}_{n}]\right]\right]=\mathbb{E}_{\mathcal{D}_{n}}\left[\mathbb{E}_{\pi_{g}}[\theta\mid\mathcal{D}_{n}]\right]=\mathbb{E}_{\pi_{g}}[\theta].$
This means that the posterior mean under $\pi_{g}$ is an unbiased estimator of the true treatment effect, even when $\mathcal{D}_{n}$ is selected adaptively through interim analyses. In other words, interim stopping does not introduce bias when the (user-specified) prior is correctly specified (i.e., $\pi=\pi_{g}$), that is, the posterior is computed using the true prior. In Bayesian terms, this is consistent with the likelihood principle: only the observed data $\mathcal{D}_{n}$ matters for inference, not the stopping rule.
However, when $\pi\neq\pi_{g}$, $\mathbb{E}_{\pi_{g}}\left[\mathbb{E}_{\mathcal{D}_{n}\mid\theta}\left[\mathbb{E}_{\pi}[\theta\mid\mathcal{D}_{n}]\right]\right]$ generally does not equal $\mathbb{E}_{\pi_{g}}[\theta]$ because the outer expectation is under $\pi_{g}$, while the inner posterior is computed using $\pi$, and $\mathbb{E}_{\pi}[\theta\mid\mathcal{D}_{n}]\neq\mathbb{E}_{\pi_{g}}[\theta\mid\mathcal{D}_{n}]$. Interim stopping selects favorable $\mathcal{D}_{n}$ with higher-than-average treatment effects, which often amplifies prior misspecification.
MSE of the posterior mean
Under the generative model $\pi_{g}$ and $f(\cdot\mid\theta)$, the MSE of $\hat{\theta}_{n}$ is given by
$\text{MSE}(\hat{\theta}_{n})=\mathbb{E}_{\theta\sim\pi_{g},\mathcal{D}_{n}\sim f(\cdot\mid\theta)}[(\hat{\theta}_{n}-\theta)^{2}]=\text{Bias}^{2}(\hat{\theta}_{n})+\text{Var}(\hat{\theta}_{n}).$
As shown above, when $\pi=\pi_{g}$, $\text{Bias}^{2}(\hat{\theta}_{n})=0$. However, early stopping on average reduces sample size (i.e., $n<N$), which in turn increase posterior variance $\text{Var}(\hat{\theta}_{n})$. Therefore, interim stopping increases MSE.
When $\pi\neq\pi_{g}$, both bias and variance contribute to an even greater increase in the MSE of the posterior mean. This inflation is further amplified by interim stopping, as stopping rules tend to preferentially select data with extreme or favorable outcomes (e.g., early signs of efficacy), thereby skewing the posterior distribution toward overly optimistic estimates. The combination of prior misspecification and selection bias from early stopping leads to substantially degraded inferential accuracy.
Coverage of credible intervals
Let $CI_{n}$ denote a $100(1-\alpha)\%$ Bayesian credible interval for $\theta$ based on $\pi(\theta)$ and $\mathcal{D}_{n}$. The coverage of this interval is defined as $\Pr_{\theta\sim\pi_{g},\mathcal{D}_{n}\sim f(\cdot\mid\theta)}(\theta\in CI_{n}).$
When $\pi=\pi_{g}$, the posterior distribution is calibrated to the true data-generating process in the sense that
$\Pr_{\theta\sim\pi_{g},\mathcal{D}_{n}\sim f(\cdot\mid\theta)}(\theta\in CI_{n})=1-\alpha,$
implying that Bayesian credible intervals maintain their nominal frequentist coverage, even under interim analyses with stopping rules (Dawid, 1982). This result follows from the law of total expectations:
$\mathbb{E}_{\theta\sim\pi_{g},\mathcal{D}_{n}\sim f(\cdot\mid\theta)}[\mathbf{1}_{\theta\in CI_{n}}]=\mathbb{E}_{\mathcal{D}_{n}}\left[\mathbb{E}_{\theta\sim\pi_{g}(\cdot\mid\mathcal{D}_{n})}[\mathbf{1}_{\theta\in CI_{n}}]\right],$
Since $\pi=\pi_{g}$, the posterior is calculated under the true prior $\pi_{g}$, and by construction of the Bayesian credible interval, the inner expectation is:
$\mathbb{E}_{\theta\sim\pi_{g}(\cdot\mid\mathcal{D}_{n})}[\mathbf{1}_{\theta\in CI_{n}}]=\Pr_{\theta\sim\pi_{g}}(\theta\in CI_{n}\mid\mathcal{D}_{n})=1-\alpha.$
Hence,
$\Pr_{\theta,\mathcal{D}_{n}}(\theta\in CI_{n})=\mathbb{E}_{\mathcal{D}_{n}}[1-\alpha]=1-\alpha.$
In contrast, when $\pi\neq\pi_{g}$, the posterior (and the resulting Bayesian credible interval) is misaligned with the data-generating process. As a result, the credible interval may under- or over-cover the true parameter, leading to deviations from nominal coverage. This is because when $\pi\neq\pi_{g}$, the posterior used to construct credible interval is misspecified, and thus the inner expectation is:
$\mathbb{E}_{\theta\sim\pi_{g}(\cdot\mid\mathcal{D}_{n})}[\mathbf{1}_{\theta\in CI_{n}}]\neq 1-\alpha.$
Therefore,
$\Pr_{\theta,\mathcal{D}_{n}}(\theta\in CI_{n})=\mathbb{E}_{\mathcal{D}_{n}}[\mathbb{E}_{\theta\sim\pi_{g}(\cdot\mid\mathcal{D}_{n})}[\mathbf{1}_{\theta\in CI_{n}}]]\neq 1-\alpha.$
This mismatch can be further amplified by interim stopping, as stopping rules tend to preferentially select data with extreme or favorable outcomes, which biases the posterior toward an overly optimistic direction. As a result, interim stopping can exacerbate the undercoverage of credible intervals when the prior is misspecified.
FDR
Recall that the false discovery rate (FDR) is defined as
$\text{FDR}=\mathbb{E}\left[\frac{V}{R\vee 1}\right].$
Under a fixed rejection threshold $C$, conducting multiple interim analyses increases the likelihood of rejecting $H_{0}$ due to random data fluctuations—particularly early in the trial. This leads to an accumulation of false rejections across interim looks, widely known as multiplicity. In contrast, the number of true rejections under the alternative hypothesis $H_{a}$ tends to grow more slowly or even plateau. This is because the statistical power is typically high by design (e.g., 0.8), and additional interim looks offer diminishing returns in detecting true effects. As a result, the expected number of false discoveries ($V$) tends to increase more rapidly than the total number of rejections ($R$), leading to inflation of the FDR (see Supplementary Materials for a proof). This inflation occurs even when the prior distribution is correctly specified (i.e., $\pi=\pi_{g}$), and may be further exacerbated under prior misspecification (i.e., $\pi\neq\pi_{g}$). A similar argument applies to the pFDR as well.
Type I error rates A and B
Bayesian interim analyses inflate type I error rate A, regardless of whether $\pi(\theta)$ is correctly or incorrectly specified. This is because at each interim analysis, there is a non-zero probability that the posterior probability exceeds the threshold, i.e., $\Pr(H_{a}\mid\mathcal{D}_{n})>C$, purely due to random fluctuations in the observed data. That is, $H_{0}$ may be falsely rejected by chance when $\theta-\theta_{0}<\delta$, resulting in inflation of the type I error rate A. This inflation is often further exacerbated under prior misspecification (i.e., $\pi\neq\pi_{g}$), as stopping rules tend to preferentially select data with extreme or favorable outcomes (e.g., indicating early efficacy), thereby biasing the posterior toward an overly optimistic direction.
Since Type I error rate B is defined based on credible intervals, one might expect that interim stopping would not inflate it when $\pi=\pi_{g}$. However, this is not the case. Because Type I error rate B is defined conditional on $H_{0}$, not all parameter values generated under $\pi_{g}$ contribute to its calculation. In effect, only the portion of $\pi_{g}$ that lies within the null parameter space (e.g., $\theta-\theta_{0}<\delta$) is relevant. This creates a mismatch between the data-generating prior $\pi_{g}$ and the prior $\pi$ used for inference, even when they are nominally the same. As a result, Type I error rate B can differ from that of a fixed design, despite using $\pi=\pi_{g}$. This deviation becomes more pronounced when $\pi\neq\pi_{g}$.
3 Simulation
3.1 Simulation setting
We conducted extensive simulation studies to evaluate the impact of repeated Bayesian interim analyses on various design metrics. The maximum sample size was set to $N=100$, and an equivalence margin of $\delta=0$ was used to reflect the setting of typical superiority trials. The data-generating prior, $\pi_{g}(\theta)$, was assumed to belong to the same probability distribution family as the user-specified prior, $\pi(\theta)$, but with potentially different hyperparameter values. For each outcome type and each scenario defined by $\pi(\theta)$, we simulated 50,000 trials. In the $k$th simulated trial, we first sampled $\theta^{k}$ from $\pi_{g}(\theta)$, then generated interim data $X^{k}_{1},X^{k}_{2},\cdots,X^{k}_{n}$ from $f(X_{i}|\theta^{k})$, e.g., $\mbox{Bernoulli}(\theta^{k})$ for binary endpoints or $N(\theta^{k},\sigma^{2})$ for normal endpoints. We implemented two interim analyses at the 40th and 70th patients, and a final analysis at the 100th patient, i.e., for n=40, 70n=40,\ 70, and $100$, the trial was stopped and $H_{0}$ rejected if $\mbox{Pr}(H_{a}|{\cal D}^{k}_{n})>C$, where ${\cal D}^{k}_{n}=\{X^{k}_{1},\cdots,X^{k}_{n}\}$.
As a benchmark, we also evaluated the same design metrics for trials conducted without interim analyses. In these trials, the decision to reject $H_{0}$ was made at the end of the trial based on the criterion $\mbox{Pr}(H_{a}|{\cal D}^{k}_{N})>C$. The probability cutoff $C$ was calibrated to control the pFDR at 5% under the scenario where $\pi(\theta)=\pi_{g}(\theta)$. This same cutoff was then applied to designs with and without interim analyses; therefore, any differences observed can be directly attributed to the inclusion of interim analyses. It is important to note that we did not re-calibrate the decision cutoff $C$ for each individual scenario. As a result, in scenarios where $\pi(\theta)\neq\pi_{g}(\theta)$, the pFDR and other metrics for the fixed design may deviate from the nominal level of 0.05. However, this does not pose a concern for our investigation, as the primary objective is to compare results with and without interim analyses to assess their impact on operating characteristics, rather than to evaluate the absolute operating characteristics of a design across different scenarios.
Below are the simulation parameters for single-arm trials; those for RCTs are provided in Supplementary Materials. For binary endpoint, we set the response rate of the (historical) control as $\theta_{0}=0.6$ and assumed the true response rate of the experimental arm generated from $\pi_{g}(\theta)=$ Beta(3, 3). We evaluated three classes of user-specified prior $\pi(\theta)$: (i) neutral priors centered at $\theta_{0}=0.6$; (ii) pessimistic priors centered at $0.5$; and (iii) optimistic priors centered at $0.7$. Within each class, we examined three priors with increasing spread (i.e., decreasing informativeness). Specifically, for $\pi(\theta)$ centered at 0.6, we considered Beta(18, 12), Beta(3.6, 2.4), and Beta(0.06, 0.04). For $\pi(\theta)$ centered at 0.5, we considered Beta(15, 15), Beta(3, 3), and Beta(0.05, 0.05). Notably, Beta(3, 3) matches $\pi_{g}(\theta)$. And for $\pi(\theta)$ centered at 0.7, we considered Beta(21, 9), Beta(4.2, 1.8), and Beta(0.07, 0.03). The probability cutoff was calibrated as previously described and set to $C=0.689$. Due to the discrete nature of the observed data, it was not always possible to find a cutoff yielding an exact 5% pFDR. This value of $C$ was selected to achieve an pFDR closest to, but not exceeding, 5%.
For the normal outcome with $\sigma^{2}=1$, we assumed $\pi_{g}(\theta)=N(0,1)$ and set $\theta_{0}=0.25$. The neutral priors $\pi(\theta)$ centered at 0.25 included $N(0.25,0.2),N(0.25,1)$, and $N(0.25,1000)$. Optimistic and pessimistic priors were centered at 0 and 0.5, respectively. The probability cutoff was calibrated and set to $C=0.4$. For the normal outcome with unknown variance, we used the conjugate prior: $\pi_{g}(\theta|\sigma^{2})=N(0,\sigma^{2}/5)$ and $\pi_{g}(\sigma^{2})=\mbox{Inv-}\chi^{2}(5,40)$, with $\theta_{0}=0.25$. Priors $\pi(\theta,\sigma^{2})$ were constructed with $\theta$ centered at 0.25 (neutral), 0 (pessimistic), and 0.5 (optimistic), as detailed in Table 4. The probability cutoff was set at $C=0.63$.
For the time-to-event outcomes, we assumed a 12-month follow-up period with patient accrual following a Poisson process at a rate of six patients per month. We set $\theta_{0}=8$ and used $\pi_{g}(\theta)=\mbox{Gamma}(12,2)$ and $\pi_{g}(\kappa)=\mbox{Gamma}(8,2)$. We evaluated three $\pi(\theta,\kappa)$ priors with $\theta$ centered at 8 (neutral), 6 (pessimistic), and 10 (optimistic), each with varying precision (see Table 5). The cutoff was calibrated and set to $C=0.61$ to control the pFDR at 5% in the absence of an interim analysis.
The performance metrics were calculated as follows. pFDR was calculated as the proportion of trials in which $H_{0}$ was true among those in which $H_{0}$ was rejected, across the 50,000 simulated trials. FDR was then obtained by multiplying the pFDR by the proportion of trials in which $H_{0}$ was rejected. Bias was determined as the average difference $\hat{\theta}^{k}-\theta^{k}$, and MSE was computed as the mean of $(\hat{\theta}^{k}-\theta^{k})^{2}$, where $\hat{\theta}^{k}$ denotes the posterior mean of $\theta$ in the $k$th simulated trial. Coverage was defined as the percentage of trials in which $\theta^{k}$ fell within the 95% posterior credible interval for $\theta$ in the $k$th trial. Type I error rate A was computed as the percentage of rejected $H_{0}$ among those in which $H_{0}$ was true, and power was computed as the percentage of rejected $H_{0}$ among those in which $H_{0}$ was false. Type I error rate B was defined as one minus the coverage probability for the one-sided posterior credible interval when $H_{0}$ was true. Additionally, we report the average sample size across all simulated trials.
3.2 Results
Table 2 presents results for single-arm trials with a binary endpoint, using the design without interim analyses as the reference. Unless otherwise specified, the following discussion focuses on comparing designs with interim analyses to those without, in order to assess the impact of conducting multiple interim looks. Our findings are consistent with the theoretical investigation in Section 2.3, indicating that interim analyses can have a substantial effect on the operating characteristics of the design.
In the ideal scenario where the user-specified prior matches the data-generating prior (i.e., $\pi(\theta)=\pi_{g}(\theta)=$Beta(3, 3)), adding interim analyses does not affect the coverage rates of either one-sided or symmetric credible intervals, or the bias. However, incorporating interim analyses does lead to inflation in pFDR, type I error A and B, power, and MSE. For example, incorporating interim analyses increased the pFDR from 0.046 to 0.095 and type I error A from 0.018 to 0.041. This differential impact across various metrics may help explain discrepancies in the literature—while some authors have claimed that interim analyses and early stopping have no effect on the operating characteristics of a design, others have argued that they do. Although both Type I error rates A and B are inflated, the inflation of the latter is substantially smaller than that of the former, suggesting that Type I error B (the Type I error rate based on credible intervals) is more robust. This robustness stems from the fact that Type I error rate B partially inherits the properties of credible intervals, which are well calibrated and remain unaffected by interim analyses when $\pi(\theta)=\pi_{g}(\theta)$.
[…]
Tables 3 and 4 present analogous results for normal endpoints with known and unknown variances, and Table 5 shows findings for time-to-event endpoints. Across all endpoint types, patterns were broadly consistent with those described above for the binary case. FDR values for these outcome types are reported in Tables S1 and S2 of the Supplementary Materials, showing similar patterns to those observed for pFDR. Results for RCTs are generally similar to those of single-arm trials; see Tables S3-S6 in Supplementary Materials for more details.
3.3 Sensitivity analyses
We examined the effect of altering the frequency of interim analyses. For binary and normal outcomes, where patient-level data are typically observed promptly, we employed a fully sequential design in place of the original two interim analyses at the 40th and 70th patients. In this design, early stopping was assessed after each patient beginning with the 20th, i.e., for $n=20,\cdots,N$, the trial was stopped and $H_{0}$ rejected if $\mbox{Pr}(H_{a}|{\cal D}^{k}_{n})>C$, where ${\cal D}^{k}_{n}=\{X^{k}_{1},\cdots,X^{k}_{n}\}$. For time-to-event endpoints, where outcomes are event-driven rather than patient-driven, interim analyses were conducted every 5 patients starting from the 20th, to allow time for meaningful accumulation of event information. Results for each outcome type under these settings are shown in Tables S7-S9 and were broadly consistent with those observed under the original design.
We examined the effect of altering the equivalence margin $\delta$. Specifically, we evaluated scenarios with positive values of $\delta$, meaning that the experimental arm must demonstrate a minimum improvement of $\delta$ (e.g., a minimal clinically important difference) over the control to be considered successful. Results for each outcome type under these settings are presented in Tables S10 - S12. The observed performance metrics followed patterns consistent with those under the original setting.
We further evaluated the performance metrics under alternative values of the probability of $H_{0}$ under $\pi_{g}(\theta)$, denoted by $\mbox{Pr}_{g}(H_{0})$. For example, for the normal outcome with known variance equal to 1 and equivalence margin $\delta=0$, the prior $\pi_{g}(\theta)=N(0,1)$ yields $\mbox{Pr}_{g}(H_{0})=0.6$. We also considered a more concentrated prior, $\pi_{g}(\theta)=N(0,0.09)$, which increases $\mbox{Pr}_{g}(H_{0})$ to 0.8. The corresponding results for continuous, binary, and time-to-event outcomes in single-arm trials, presented in Tables S13 - S15 of the Supplementary Materials, exhibited patterns consistent with those observed under the original prior specification.
For the normal outcome with unknown variance, the scale parameter $\sigma_{0}^{2}$ in $\pi(\sigma^{2})$ was originally set to 40, matching that in $\pi_{g}(\sigma^{2})$. In sensitivity analyses, we varied $\sigma^{2}_{0}$ to 20 and 60. The resulting metrics, shown in Tables S16 and S17, remained consistent with those from the main setting, further supporting the robustness of the findings.
4 Conclusion
We have presented a comprehensive investigation into the effects of repeated Bayesian interim analyses on various design metrics, including bias, MSE, the coverage probability of credible intervals, FDR, Type I error rates, and power. Our theoretical results establish, under general conditions and without reliance on any specific probability model, that repeated Bayesian interim analyses are not consequence free: they can systematically impact these inferential quantities. For some metrics (e.g., FDR), distortions arise even when the user-specified prior exactly matches the data-generating prior, and can be further exacerbated under prior misspecification. These theoretical insights are validated by extensive simulation studies, which confirm that repeated Bayesian interim analyses can affect operating characteristics of trials, particularly when vague or noninformative priors are used. This is of great concern as, in practice, noninformative or vague priors are most likely to be used.
These findings highlight the importance of accounting for the cumulative impact of repeated Bayesian interim decisions on the posterior distribution and subsequent inference. To safeguard against erroneous inference, researchers should consider strategies such as calibrating the decision threshold $C$ to account for the impact of interim analyses. This adjustment parallels multiplicity control methods in the frequentist paradigm, such as adopting stricter significance thresholds. An alternative approach is to specify a conservative, informative user-specified prior $\pi(\theta)$ that offsets the impact of interim analyses without requiring adjustment to $C$. However, this approach may be challenging to interpret. Equally critical is the need for thoughtful prior elicitation and comprehensive sensitivity analyses to evaluate the impact of prior misspecification. While Bayesian methods offer powerful tools for flexible and adaptive decision-making, their reliability hinges on a clear understanding of how repeated analyses interact with prior assumptions and inferential goals.
Disclaimer
This article reflects the views of the author and should not be construed to represent FDA’s views or policies.
Data Availability
Data sharing is not applicable in this paper as all data in this paper are computer simulated.